OpenStudy (anonymous):

subtract simplify by collecting like radical terms subtract simplify by collecting like radical terms @Mathematics

OpenStudy (anonymous):

\[7\sqrt{45}-3\sqrt{5}\]

OpenStudy (hoblos):

\[18\sqrt{5}\]